A body A is projected upwards with a velocity 98m/s. The second body B is projected with the same initial velocity but after 4s. Both the bodies will meet after

Grade:11

4 Answers

Dhulipala Srikanth
8 years ago
78.4
Shubham Swastik Behera
8 years ago
78.4 is the difference of their heights after 10 seconds.
The answer is 12 seconds.
Vishwa
6 years ago
The total time of flight of the body is 10s... So naturally the instant at which they meet will be more than 10s...hence 12sSecondly.... Use s=ut+1/2at2Second body will be t-4i.e ut+1/2at2=u(t-4)+1/2a(t-4)2
sana
4 years ago
since both bodies cover equal distance, h1 = h2
here g is negative as the body is projected upwards
therefore, ut -1/2gt^2 = u(t-4) -1/2g(t-4)^2
              ut-1/2gt^2 =  ut – 4u -1/2g(t^2 -8t+16)
               ut-1/2gt^2 =  ut – 4u -1/2gt^2 + 4gt – 8g
                              = -4u – 8g + 4gt
              4u + 8g     =4gt
substituting values, 4*98+8*9.8 = 4*9.8*t
                           4*98+8*9.8/4*9.8 = t
                          4*98/4*9.8 +8*9.8/4*9.8 = t
                              10+ 2 = t
                                   t= 12seconds